# Error Bound Analysis for the Regularized Loss of Deep Linear Neural Networks

### v1

@@ Abstract
The optimization foundations of deep linear networks have received significant attention lately. However, due to the non-convexity and hierarchical structure, analyzing the regularized loss of deep linear networks remains a challenging task. In this work, we study the local geometric landscape of the regularized squared loss of deep linear networks, providing a deeper understanding of its optimization properties. Specifically, we characterize the critical point set and establish an error-bound property for all critical points under mild conditions. Notably, we identify the sufficient and necessary conditions under which the error bound holds. To support our theoretical findings, we conduct numerical experiments demonstrating that gradient descent exhibits linear convergence when optimizing the regularized loss of deep linear networks.
@@ 1 Introduction
Deep learning has been widely used in various fields, including computer vision [16], natural language processing [39], and healthcare [12], due to its exceptional empirical performance. Optimization is a key component of deep learning, playing a pivotal role in formulating learning objectives, training neural networks, and improving model generalization. In general, optimization problems arising in deep learning are highly non-convex and difficult to analyze due to the inherent nonlinearity and hierarchical structures of deep neural networks. Even in the context of linear neural networks, which are the most basic form of neural networks, our theoretical understanding remains far from complete and systematic, especially concerning the optimization properties. This motivates us to study the following problem based on deep linear networks:
$\min_{\bm{W}}\ \left\|\bm{W}_{L}\ldots\bm{W}_{1}\bm{X}-\bm{Y}\right\|_{F}^{2}+\sum_{l=1}^{L}\lambda_{l}\|\bm{W}_{l}\|_{F}^{2},$ (1)
@@
where $(\bm{X},\bm{Y})\in\mathbb{R}^{d_{0}\times N}\times\mathbb{R}^{d_{L}\times N}$ denotes the data input, $\bm{W}_{l}\in\mathbb{R}^{d_{l}\times d_{l-1}}$ denotes the $l$-th weight matrix for each $l=1,\dots,L$, $\bm{W}=\{\bm{W}_{l}\}_{l=1}^{L}$ denotes the collection of all weight matrices, and $\lambda_{l}>0$ for all $l\in[L]$ are regularization parameters. Notably, such a problem captures a wide range of deep learning problems arising in applications, including deep matrix factorization [2, 8, 38], neural collapse [14, 44, 50], and low-rank adaption (LoRA) of large language models [18, 45], to name just a few. Moreover, studying linear networks provides a valuable starting point for gaining insights into nonlinear networks, as they exhibit similar learning behaviors and phenomena to their nonlinear counterparts while maintaining a simpler structure [20, 35].
@@
In practice, (stochastic) gradient descent (GD) and its variants are among the most widely used first-order methods for deep learning [25]. Over the past few years, substantial progress has been made in studying the convergence behavior of GD for solving Problem (1). In the literature, a large amount of work has been dedicated to studying an unregularized counterpart of Problem (1), i.e., $\min\left\|\bm{W}_{L}\ldots\bm{W}_{1}\bm{X}-\bm{Y}\right\|_{F}^{2}$, based on a gradient dynamic analysis. A noteworthy assumption in most of these studies is that the data is whitened, i.e., $\bm{X}\bm{X}^{T}=\bm{I}_{d}$, when analyzing the convergence rate of GD. In general, this assumption simplifies the analysis by ensuring that the input samples are uncorrelated and have unit variance. Now, suppose that the data is whitened. Bartlett et al. [5] showed that GD with the identity initialization converges to an $\epsilon$-optimal solution within a polynomial number of iterations. Later, Arora et al. [3] further improved the convergence result, showing that GD converges linearly to a global optimum when $\min\{d_{1},\dots,d_{L-1}\}\geq\min\{d_{0},d_{L}\}$, the initial weight matrices are approximately balanced, and the initial loss is smaller than a threshold. Meanwhile, Wu et al. [43] showed that gradient descent with zero-asymmetric initialization avoids saddle points and converges to an $\epsilon$-optimal solution in $O(L^{3}\log(1/\epsilon))$ iterations. Other works also proved similar global convergence and convergence rate results of GD under different assumptions; see, e.g., [4, 6, 19, 32, 43, 49]. Despite these inspiring results, the existing analyses suffer from three notable limitations. First, they are highly specific to a particular problem, relying on the analysis of gradient dynamics under carefully designed weight initialization schemes. This raises questions about the generalizability of these analyses to GD or other first-order optimization methods with different initialization schemes. Second, regularization is commonly used during the training of neural networks to prevent overfitting, improve generalization, and accelerate convergence [26, 48]. However, the existing gradient dynamics analyses mainly focus on unregularized deep neural networks and cannot be directly applied to Problem (1), even when the data is whitened. Finally, the existing analyses only apply to analyze the convergence to global optimal solutions. However, Problem (1) and its regularized counterpart may have local minimizers, to which first-order methods, such as GD, are likely to converge. To the best of our knowledge, the convergence behavior of first-order methods when they approach a critical point—whether a global minimum, a local minimum, or even a saddle point—remains an open question in the literature. To sum up, it remains an unsolved challenge in the literature to develop a unified framework to analyze the convergence behavior of first-order methods to critical points when solving Problem (1).
To address the above challenge, a promising approach is to study the local geometric structure of Problem (1) associated with its objective function, such as the error bound condition [29, 34, 51], the Polyak-Łojasiewicz (PŁ) inequality [33], and quadratic growth [9]. When the data is whitened, as is commonly assumed in the literature (see the review above), we let $\bm{Y}:=\bm{Y}\bm{X}^{T}$ with a slight abuse of notation. Under this notation, Problem (1) reduces to
$\min_{\bm{W}}F(\bm{W}):=\left\|\bm{W}_{L}\ldots\bm{W}_{1}-\bm{Y}\right\|_{F}^{2}+\sum_{l=1}^{L}\lambda_{l}\|\bm{W}_{l}\|_{F}^{2}.$ (2)
@@
Notably, both deep matrix factorization [2, 8, 38] and the neural collapse problem [14, 50] are special instances of the above problem. In this work, we mainly focus on the error bound condition associated with Problem (2). Formally, let $\cal W$ denote the set of critical points to Problem (2). We say that it possesses an error bound for $\cal W$ if there exist constants $\epsilon,\kappa>0$ such that for all $\bm{W}$ satisfying $\mathrm{dist}(\bm{W},\mathcal{W})\leq\epsilon$,
$\mathrm{dist}(\bm{W},\mathcal{W})\leq\kappa\|\nabla F(\bm{W})\|_{F},$ (3)
@@
where $\mathrm{dist}(\bm{W},\mathcal{W}):=\min\{\|\bm{W}-\bm{X}\|_{F}:\bm{X}\in\mathcal{W}\}$ denotes the distance from $\bm{W}$ to the critical point set $\cal W$. Intuitively, the error bound inequality (3) requires the distance from a point to the set of critical points to be bounded by its gradient norm. A unified framework leveraging this condition, along with some algorithm-dependent conditions, has been widely studied to analyze linear convergence of first-order methods [9, 51] or superlinear convergence of second-order methods [47] in convex optimization. Recently, Liao et al. [29] showed that even if the objective function is non-convex but smooth and satisfies the error bound, GD converges linearly to the solution set. Moreover, Liao et al. [29], Rebjock and Boumal [34] demonstrated that the error bound is equivalent to other regularity conditions, such as the PŁ inequality [33] and quadratic growth [9] under mild conditions. These regularity conditions are widely used in the literature to prove linear convergence of GD for optimizing non-convex problems [13]. Notably, this type of convergence analysis framework is not limited to studying convergence to global optimal solutions. Instead, it can also be applied to local minima or even saddle points, as long as the error bound holds for the targeted solution sets. This makes such a framework particularly effective for analyzing non-convex optimization problems, where the landscape often contains various types of critical points.
@@
However, powerful as this approach may seem, a critical challenge is to prove the error bound inequality (3) for Problem (2). As far as we know, proving the error bound or other regularity conditions for deep neural networks remains relatively underexplored. Recently, Wang et al. [40] made progress in this direction by establishing the error bound for the set of global optimal solutions to a special instance of Problem (2): $\min\|\bm{W}_{2}\bm{W}_{1}-\bm{Y}\|_{F}^{2}+\lambda_{1}\|\bm{W}_{1}\|_{F}^{2}+\lambda_{2}\|\bm{W}_{2}\|_{F}^{2}$, where $\lambda_{1},\lambda_{2}>0$ are regularization parameters and $\bm{Y}$ is a membership matrix. However, this analysis is limited to global optimal solutions of 2-layer neural networks and a specific data matrix $\bm{Y}$. Extending these results to deeper networks, more general data, and broader sets of critical points remains an open problem.
@@
In view of the above discussion, our goal in this paper is to establish an error bound for the set of critical points of Problem (2). Our main contributions are twofold. First, we explicitly characterize the critical point set of Problem (2) for general data $\bm{Y}$ (see 1), despite the non-convexity and hierarchical structure of the problem. This characterization serves as a foundation for establishing the error bound for Problem (2). Second, leveraging this explicit characterization, we show that all critical points of Problem (2) satisfy the error bound (see 2) under mild conditions. Notably, we identify the sufficient and necessary conditions on the relationship between the regularization parameters and the spectrum of the input data that ensure the error bound holds. Such a result is significant, as it expands the currently limited repertoire of non-convex problems for which the local loss geometry is well understood. Moreover, it is important to note that our work develops new techniques in the process of establishing the error bound to handle the repeated singular values in $\bm{Y}$ and the complicated structure of the critical point set. These techniques could be of independent interest. The established error bound can be used to establish other regularity conditions, such as the PŁ inequality and quadratic growth. Building on the error bound of Problem (2), we demonstrate that first-order methods can achieve linear convergence to a critical point of Problem (2), provided that they satisfy certain algorithm-dependent properties. We conduct numerical experiments in deep linear networks and in more general settings and observe that gradient descent converges linearly to critical points. These results strongly support our theoretical findings.
Organization.
The rest of this paper is organized as follows. In Section 2, we present the main results of this paper. We introduce key lemmas and propositions without proofs and prove the main results in Section 3, and provide the detailed proofs for these lemmas and propositions in Section 4. We report experimental results in Section 5 and conclude the paper in Section 6.
Notation.
@@
Given an integer $n$, we denote by $[n]$ the set $\{1,\dots,n\}$. Given a vector $\bm{a}$, let $\|\bm{a}\|$ denote the Euclidean norm of $\bm{a}$, $a_{i}$ the $i$-th entry, and $\mathrm{diag}(\bm{a})$ the diagonal matrix with $\bm{a}$ as its diagonal. Unless specified otherwise, all vectors in this paper are column vectors. Given a matrix $\bm{A}\in\mathbb{R}^{N\times d}$, let $\|\bm{A}\|$ denote the spectral norm of $\bm{A}$, $\|\bm{A}\|_{F}$ denote the Frobenius norm, $a_{ij}$ denote the $(i,j)$-th element, and $\sigma_{i}(\bm{A})$ denote the $i$-th largest singular value. We use $\bm{0}_{m\times n}$ to denote $m\times n$ all-zero matrix, $\bm{0}_{m}$ to denote $m\times m$ all-zero matrix, and simply write $\bm{0}$ when its dimension can be inferred from the context. We use $\mathcal{O}^{n\times d}$ to denote the set of all $n\times d$ matrices that have orthonormal columns (in particular, we use $\mathcal{O}^{n}$ to denote the set of all $d\times d$ orthogonal matrices); $\mathcal{P}^{n}$ to denote the set of all $n\times n$ permutation matrices; $\mathrm{BlkDiag}(\bm{X}_{1},\dots,\bm{X}_{n})$ to denote the block diagonal matrix whose diagonal blocks are $\bm{X}_{1},\dots,\bm{X}_{n}$. Given a matrix $\bm{X}\in\mathbb{R}^{m\times n}$ and a non-empty closed set $\mathcal{X}\subseteq\mathbb{R}^{m\times n}$, we use $\mathrm{dist}(\bm{X},\mathcal{X})$ to denote the Euclidean distance from $\bm{X}$ to $\cal X$; the distance between $\mathcal{X}$ and another non-empty closed set $\mathcal{Y}$ is defined as $\mathrm{dist}(\mathcal{X},\mathcal{Y})=\min_{\bm{X}\in\mathcal{X},\bm{Y}\in\mathcal{Y}}\|\bm{X}-\bm{Y}\|_{F}$. Given weight matrices $\bm{W}_{1},\dots,\bm{W}_{L}$, let $\bm{W}_{i:1}:=\bm{W}_{i}\bm{W}_{i-1}\ldots\bm{W}_{1}$ for each $i=2,\dots,L$ and $\bm{W}_{L:i}:=\bm{W}_{L}\bm{W}_{L-1}\ldots\bm{W}_{i}$ for each $i=1,\dots,L-1$. In particular, we define $\bm{W}_{0:1}:=\bm{I}$ and $\bm{W}_{L:L+1}:=\bm{I}$. For all $i\geq j+1$, let $\bm{W}_{i:j}=\bm{W}_{i}\bm{W}_{i-1}\ldots\bm{W}_{j}$ and $\bm{W}_{i:j}=\bm{W}_{i}$ when $i=j$. All other notation is standard.

### v3

@@ Abstract
The optimization foundations of deep linear networks have recently received significant attention. However, due to their inherent non-convexity and hierarchical structure, analyzing the loss functions of deep linear networks remains a challenging task. In this work, we study the local geometric landscape of the regularized squared loss of deep linear networks around each critical point. Specifically, we derive a closed-form characterization of the critical point set and establish an error bound for the regularized loss under mild conditions on network width and regularization parameters. Notably, this error bound quantifies the distance from a point to the critical point set in terms of the current gradient norm, which can be used to derive linear convergence of first-order methods. To support our theoretical findings, we conduct numerical experiments and demonstrate that gradient descent converges linearly to a critical point when optimizing the regularized loss of deep linear networks.
@@ 1 Introduction
Deep learning has achieved great empirical success in various fields, including computer vision [17], natural language processing [41], and healthcare [13]. Optimization is a key component of deep learning, playing a pivotal role in formulating learning objectives and training neural networks. In general, optimization problems arising in deep learning are highly non-convex and difficult to analyze due to the inherent nonlinearity and hierarchical structures of deep neural networks. Even in the context of linear neural networks, which are the most basic form of neural networks, our theoretical understanding remains far from complete and systematic. In this work, we study the following problem based on deep linear networks:
$\min_{\bm{W}}\ F(\bm{W}):=\left\|\bm{W}_{L}\cdots\bm{W}_{1}-\bm{Y}\right\|_{F}^{2}+\sum_{l=1}^{L}\lambda_{l}\|\bm{W}_{l}\|_{F}^{2},$ (1)
@@
where $L\geq 2$ denotes the number of layers, $\bm{Y}\in\mathbb{R}^{d_{L}\times d_{0}}$ denotes the data input, $\bm{W}_{l}\in\mathbb{R}^{d_{l}\times d_{l-1}}$ denotes the $l$-th weight matrix, $\bm{W}=\{\bm{W}_{l}\}_{l=1}^{L}$ denotes the collection of all weight matrices, and $\lambda_{l}>0$ for all $l$ are regularization parameters. Notably, such a problem captures a wide range of deep learning problems, including deep matrix factorization [2, 8, 40], neural collapse [15, 50], and low-rank adaptation of learning models [19, 45].
In the literature, a large amount of work has been dedicated to studying the following problem:
$\min_{\bm{W}}\left\|\bm{W}_{L}\cdots\bm{W}_{1}\bm{X}-\bm{Z}\right\|_{F}^{2},$
where $(\bm{X},\bm{Z})\in\mathbb{R}^{d_{0}\times N}\times\mathbb{R}^{d_{L}\times N}$ denotes the data input. A common assumption in many theoretical studies (see, e.g., [1, 5, 9, 44]) is that the data is whitened, i.e., $\bm{X}\bm{X}^{T}=\bm{I}_{d}$. Under this assumption, one can verify that $\min_{\bm{W}}\left\|\bm{W}_{L}\cdots\bm{W}_{1}\bm{X}-\bm{Z}\right\|_{F}^{2}$ is equivalent to $\min_{\bm{W}}\left\|\bm{W}_{L}\cdots\bm{W}_{1}-\bm{Z}\bm{X}^{T}\right\|_{F}^{2}$. Obviously, this formulation is an unregularized counterpart of Problem (1). We should point out that regularization is commonly used during the training of neural networks to prevent overfitting, improve generalization, and accelerate convergence [25, 48]. However, the existing theoretical studies mainly focus on unregularized deep neural networks and cannot be directly applied to Problem (1), even when the data is whitened; see, e.g., [1, 4, 5, 7, 20, 32, 44].
@@
In practice, (stochastic) gradient descent (GD) and its variants are among the most widely used first-order methods for deep learning [24]. Over the past few years, substantial progress has been made in studying the convergence behavior of GD for optimizing deep linear networks, particularly in the context of the unregularized problem $\min_{\bm{W}}\left\|\bm{W}_{L}\cdots\bm{W}_{1}-\bm{Y}\right\|_{F}^{2}$. For example, Bartlett et al. [5] showed that GD with the identity initialization converges to an $\epsilon$-optimal solution within a polynomial number of iterations. Later, Arora et al. [1] further improved the convergence result, showing that GD converges linearly to a global optimum when $\min\{d_{1},\dots,d_{L-1}\}\geq\min\{d_{0},d_{L}\}$, the initial weight matrices are approximately balanced, and the initial loss is smaller than a threshold. Meanwhile, Wu et al. [44] showed that gradient descent with zero-asymmetric initialization avoids saddle points and converges to an $\epsilon$-optimal solution in $O(L^{3}\log(1/\epsilon))$ iterations. Other works also proved similar global convergence and convergence rate results of GD under different assumptions; see, e.g., [4, 7, 20, 32, 44]. Despite these inspiring results, the existing analyses suffer from some notable limitations. First, they are highly specific to a particular problem, relying on the analysis of gradient dynamics under carefully designed weight initialization schemes. In addition, the existing analyses only apply to analyze the convergence to global optimal solutions. However, Problem (1) and its unregularized counterpart may have other local solutions, to which first-order methods, such as GD, are likely to converge. To the best of our knowledge, the convergence behavior of first-order methods when they approach a critical point—whether a global minimum, a local minimum, or even a saddle point—remains an open question in the literature. To sum up, it remains an unsolved challenge in the literature to develop a unified framework to analyze the convergence behavior of first-order methods when solving Problem (1).
@@
To address the above challenge, a promising approach is to study the local geometric structure of Problem (1) associated with its objective function, such as the error bound condition [29, 36, 51], the Polyak-Łojasiewicz (PŁ) inequality [34], and quadratic growth [10]. In this work, we mainly focus on the error bound condition associated with Problem (1). Formally, let $\mathcal{W}_{F}$ denote the set of critical points of Problem (1). We say that it possesses an error bound for $\mathcal{W}_{F}$ if there exist constants $\epsilon,\kappa>0$ such that for all $\bm{W}$ satisfying $\mathrm{dist}(\bm{W},\mathcal{W}_{F})\leq\epsilon$, we have
$\mathrm{dist}(\bm{W},\mathcal{W}_{F})\leq\kappa\|\nabla F(\bm{W})\|_{F},$ (2)
@@
where $\mathrm{dist}(\bm{W},\mathcal{W}_{F}):=\min\{\|\bm{W}-\bm{V}\|_{F}:\bm{V}\in\mathcal{W}_{F}\}$ denotes the distance from $\bm{W}$ to the critical point set $\mathcal{W}_{F}$. Intuitively, the error bound inequality (2) requires the distance from a point to the set of critical points to be bounded by its gradient norm. A unified framework leveraging this condition, along with some algorithm-dependent conditions, has been widely studied to analyze linear convergence of first-order methods [10, 51] or superlinear convergence of second-order methods [47] in convex optimization. Luo and Tseng [31] showed that even if the objective function is non-convex but smooth and satisfies the error bound, (projected) GD converges linearly to the solution set. Moreover, Li and Pong [28], Liao et al. [29], Rebjock and Boumal [36] demonstrated that the error bound is equivalent to other regularity conditions, such as the PŁ inequality [34] and quadratic growth [10] under mild conditions. These regularity conditions are widely used in the literature to prove linear convergence of GD for optimizing non-convex problems [34, 31, 3, 14]. Notably, this type of convergence analysis framework is not limited to studying convergence to global optimal solutions. Instead, it can also be applied to local minima or even saddle points, as long as the error bound holds for the targeted solution sets. This makes such a framework particularly effective for analyzing non-convex optimization problems, whose landscape often contains various types of critical points.
@@
However, powerful as this approach may seem, proving error bounds for specific problems, particularly non-convex ones, is technically difficult, as it often requires a delicate analysis [51, 21, 22, 30, 43, 33]. As far as we know, proving the error bound or other regularity conditions for deep neural networks remains relatively underexplored. Recently, Wang et al. [42] made progress in this direction by establishing the error bound for the set of global optimal solutions to a special instance of Problem (1): $\min\|\bm{W}_{2}\bm{W}_{1}-\bm{Y}\|_{F}^{2}+\lambda_{1}\|\bm{W}_{1}\|_{F}^{2}+\lambda_{2}\|\bm{W}_{2}\|_{F}^{2}$, where $\lambda_{1},\lambda_{2}>0$ are regularization parameters and $\bm{Y}$ is a membership matrix. Moreover, Zhang et al. [49] and Tao et al. [39] extended the analysis to the setting of general data input $\bm{Y}$. However, these analyses are limited to global optimal solutions of 2-layer linear networks. Extending these results to deeper networks and critical points remains an open problem.
@@
In view of the above discussion, our goal in this paper is to establish an error bound for the set of critical points of Problem (1). Our main contributions are twofold. First, we explicitly characterize the critical point set of Problem (1) for general data $\bm{Y}$ (see 1), despite the non-convexity and hierarchical structure of the problem. This characterization serves as a foundation for establishing the error bound for Problem (1). Second, leveraging this explicit characterization, we show that all critical points of Problem (1) satisfy the error bound (see 2) under mild conditions on the network width and regularization parameters. More specifically, we identify the sufficient and necessary conditions on the relationship between the regularization parameters and the spectrum of the input data that ensure the error bound holds. Such a result is significant, as it expands the currently limited repertoire of non-convex problems for which the local loss geometry is well understood. Moreover, it is important to note that our work develops new techniques in the process of establishing the error bound to handle the repeated singular values in $\bm{Y}$ and the complicated structure of the critical point set. These techniques could be of independent interest. The established error bound can be used to establish other regularity conditions, such as the PŁ inequality and quadratic growth. Building on the error bound of Problem (1), we demonstrate that first-order methods can achieve linear convergence to a critical point of Problem (1), provided that they satisfy certain algorithm-dependent properties. Finally, we conduct numerical experiments in deep linear networks and in more general settings and observe that gradient descent converges linearly to critical points. These results strongly support our theoretical findings.
@@
Notation. Given a positive integer $n$, let $n]:=\{1,\dots,n\}$. Given a set of numbers $\{\lambda_{1},\dots,\lambda_{n}\}$, we define $\lambda_{\min}:=\min\{\lambda_{i}:i\in[n]\}$ and $\lambda_{\max}:=\max\{\lambda_{i}:i\in[n]\}$. Given a vector $\bm{a}$, let $\|\bm{a}\|$ denote its Euclidean norm, $\|\bm{a}\|_{0}$ its $\ell_{0}$ norm (i.e., number of non-zero entries), $a_{i}$ the $i$-th entry, and $\mathrm{diag}(\bm{a})$ the diagonal matrix with $\bm{a}$ on the diagonal. Given a matrix $\bm{A}$, let $\|\bm{A}\|$ denote the spectral norm, $\|\bm{A}\|_{F}$ the Frobenius norm, $a_{ij}$ the $(i,j)$-th element, and $\sigma_{i}(\bm{A})$ the $i$-th largest singular value. We use $\bm{0}_{m\times n}$ to denote an all-zero matrix of size $m\times n$, $\bm{0}_{m}$ to denote all-zero vector of dimension $m$, and simply write $\bm{0}$ when the dimension can be inferred from the context. We denote the $i$-th standard basis vector by $\bm{e}_{i}$. We simply write $\nabla_{\bm{W}_{l}}F(\bm{W})$ as $\nabla_{l}F(\bm{W})$.
We use $\mathcal{O}^{n\times d}$ to denote the set of all $n\times d$ matrices that have orthonormal columns (in particular, $\mathcal{O}^{n}$ for $n\times n$ orthogonal matrices); $\mathcal{P}^{n}$ to denote the set of all $n\times n$ permutation matrices; $\mathrm{BlkD}(\bm{X}_{1},\dots,\bm{X}_{n})$ to denote the block diagonal matrix whose diagonal blocks are $\bm{X}_{1},\dots,\bm{X}_{n}$. Given a matrix $\bm{X}\in\mathbb{R}^{m\times n}$ and a non-empty closed set $\mathcal{X}\subseteq\mathbb{R}^{m\times n}$, the Euclidean distance from $\bm{X}$ to $\cal X$ is defined as $\mathrm{dist}(\bm{X},\mathcal{X})=\min_{\bm{Z}\in\mathcal{X}}\|\bm{X}-\bm{Z}\|_{F}$; the distance between $\mathcal{X}$ and a closed set $\mathcal{Y}\neq\emptyset$ is defined as $\mathrm{dist}(\mathcal{X},\mathcal{Y})=\min_{\bm{X}\in\mathcal{X},\bm{Y}\in\mathcal{Y}}\|\bm{X}-\bm{Y}\|_{F}$. Let $\pi:[d]\rightarrow[d]$ denote a permutation of the elements in $[d]$ with the corresponding permuation matrix $\bm{\Pi}\in\mathcal{P}^{d}$ satisfying $\bm{\Pi}(i,j)=1$ if $j=\pi(i)$ and $\bm{\Pi}(i,j)=0$ otherwise for each $i\in[d]$. Given weight matrices $\{\bm{W}_{l}\}_{l=1}^{L}$ with $\bm{W}_{l}\in\mathbb{R}^{d_{l}\times d_{l-1}}$ for each $l\in[L]$, let $\bm{W}_{i:1}:=\bm{W}_{i}\bm{W}_{i-1}\cdots\bm{W}_{1}$ for each $i=2,\dots,L$ and $\bm{W}_{L:i}:=\bm{W}_{L}\bm{W}_{L-1}\cdots\bm{W}_{i}$ for each $i=1,\dots,L-1$. In particular, we define $\bm{W}_{0:1}:=\bm{I}$ and $\bm{W}_{L:L+1}:=\bm{I}$. For all $i\geq j+1$, let $\bm{W}_{i:j}=\bm{W}_{i}\bm{W}_{i-1}\cdots\bm{W}_{j}$ and $\bm{W}_{i:j}=\bm{W}_{i}$ when $i=j$. All other notations are standard.
Organization. In Section 2, we present the main results of this paper. In Section 3, we present the proofs of the main results. We report experimental results in Section 4 and conclude the paper in Section 5.